# A Comprehensive Study on Large Language Models for Mutation Testing

### v4

@@ Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we conduct a comprehensive empirical study evaluating BugFarm and LLMorpheus (the two state-of-the-art LLM-based approaches), alongside seven LLMs using our newly designed prompt, including both leading open- and closed-source models, on 851 real bugs from two Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection. That is, 87.98% (for LLMs) vs. 41.64% (for rule-based); an increase of percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points, respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
@@ 1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In the most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Harman et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021b; Petrović and Ivanković, 2018) reported on results from the deployment of mutation testing at their companies. In particular, results recently reported by Meta (Harman et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025). These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
@@
Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
@@
Generating valid mutations.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutations, to judge whether the mutants are killed. Ensuring executable mutations is challenging because mutations must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
@@
Generating diverse and killable mutations: Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
@@
This paper examines how well Large-Language Models (LLMs) address these challenges, in comparison to more traditional rule-based approaches. We conducted a comprehensive empirical study which studies BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutation generation approaches), alongside seven LLMs adopting our newly designed prompt, covering both leading open- and closed-source models, and four existing mutation testing techniques (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)) with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our primary experiments required the generation of 666,148 mutations.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies, turbulence on prompt, and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), DeepSeek-V3 (Liu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. We also explore the potential of using LLMs for generating high-order mutations. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
@@
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This results in a higher fault detection rate compared to traditional approaches. For instance, the mutants generated by DeepSeek-V3-761b and GPT-4o-Mini achieve fault detection rates of 91.1 % and 90.8%, respectively, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that the DeepSeek models exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The existing LLM-based approach, BugFarm, introduces 31 node types, higher than the traditional approaches as well. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also show that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have an AST edit distance of one, whereas for GPT-4o, only 50% of its mutations have the same distance.
@@
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
@@
We analyze the alive (i.e., survive) mutations of each approach, which are important to strengthen the test case. For one alive non-equivalent mutation, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection. We analyze the distribution of alive mutations across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutations, which consequently remain alive, providing more valuable guidance for enhancing unit tests.
Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
@@
We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our study not only covers other existing LLM-based mutation generation approaches, but also introduces a new effective prompt template. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
We construct a high-quality dataset of Java mutations, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
@@ 2. Study Design
In this section, we introduce how our study is designed and conducted.
@@ 2.1. Overview and Research Questions
This paper intends to evaluate existing mutation generation approaches by answering the following research questions.
@@
RQ1 (Effectiveness) How well do mutations generated by LLMs mimic real bugs compared to existing methods? This research question investigates the effectiveness of LLM-generated mutations in representing real bugs. Specifically, it evaluates whether these mutations exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity. This RQ is divided into the following sub-RQs:
@@
RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutations?
@@
RQ1.2 (Coupling) To what extent do LLM-generated mutations couple with real bugs?
RQ1.3 (Diversity) How diverse are LLM-generated mutations compared to traditional approaches?
@@
RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity? This research question explores the validity of mutations generated by LLMs compared to traditional approaches. Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutations. By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutations that contribute to practical mutation testing. This RQ is divided into the following sub-RQs:
@@
RQ2.1 (Compilability) What percentage of generated mutations are compilable across different approaches?
@@
RQ2.2 (Duplication) What percentage of the generated mutations are duplicates of the original code across different approaches?
@@
RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
@@
RQ3 (Efficiency) What are the general costs of different mutation generation approaches? This research question investigates the general costs of different mutation generation approaches. We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutation generation approaches.
@@
RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches? This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutations generated by different approaches. Understanding these errors can guide mutation generation approaches to enhance their practical usability.
@@
RQ5 (Alive Mutations) What are the distributions of alive mutations across different approaches? This RQ aims to investigate the underlying reasons why mutations remain alive, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests. By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutations and their potential value for guiding test suite improvement.
2.2. Studied Mutation Generation Approaches
@@
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches are for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. The details of the settings for each generation approach are illustrated in Section 2.5. Table 1 summarizes the mutation generation approaches covered by this study.
@@
| BERT | Mutation generation with the pre-trained model CodeBERT. | |
@@
| LLMph (GPT-4o) | The existing LLM-based mutation generation approach LLMorpheus with GPT-4o. | |
@@
| LLMut (DS-671b) | Our approach with DeepSeek-V3-761b, whose base model is DeepSeek, released in 2024/12. |

### v5

@@ Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we conduct a comprehensive empirical study evaluating both a wide range of traditional approaches and LLM-based approaches. Particularly, we evaluate BugFarm, LLMorpheus, and our newly designed prompt for mutant generation. The experiments cover both leading open- and closed-source LLMs, on 851 real bugs from two Java real-world bug benchmarks (i.e., 605 bugs from 12 projects of Defects4J 2.0 and 246 bugs of ConDefects). Our results reveal that, compared to existing traditional approaches, LLMs generate more diverse mutants that are behaviorally closer to real bugs and, most importantly, achieve a 1.75 improvement in real bug detection. Specifically, LLM-based approaches reach a detection rate of 76.47%, compared to 44.15% for rule-based techniques, representing an absolute gain of percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points, respectively. These findings provide actionable insights for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
@@ 1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019; Potter et al., 2025). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In the most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Harman et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021b; Petrović and Ivanković, 2018) reported on results from the deployment of mutation testing at their companies. In particular, results recently reported by Meta (Harman et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutants by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025). These methods aim to introduce mutants that better resemble real faults, potentially increasing their relevance for testing.
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study exists. Without such a study, there is no reliable foundation for confidently deploying the technique at an industrial scale. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutant generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
@@
Generating effective mutants: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
@@
Generating valid mutants.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutants, to judge whether the mutants are killed. Ensuring executable mutants is challenging because mutants must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
@@
Generating diverse and killable mutants: Determining whether a mutant is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutants need to be diverse and cover the entire spectrum of code and program behaviors.
@@
This paper examines how well Large Language Models (LLMs) address these challenges compared to more traditional rule-based approaches. We conducted a comprehensive empirical study that studies BugFarm (Ibrahimzada et al., 2025) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutant generation approaches). Moreover, we propose a novel prompt template for mutant generation, which is equipped with few-shot examples (referred to as LLMut for ease of presentation). All three methods are evaluated on two representative LLMs, GPT-4o and DeepSeek-V3-671b (Liu et al., 2024) (DS-671b), denoted as BugFarm (GPT-4o), BugFarm (DS-671b), LLMorpheus (GPT-4o), LLMorpheus (DS-671b), LLMut (GPT-4o), and LLMut (DS-671b), respectively. Furthermore, LLMut is additionally evaluated on five other widely used models: GPT-3.5-Turbo (Achiam et al., 2023) (GPT-3.5), GPT-4o-Mini (GPT-4o-M), DeepSeek-Coder-V2-236b (Zhu et al., 2024) (DS-236b), StarChat--16b (Li et al., 2023a) (SC-16b), and CodeLlama-Instruct-13b (Roziere et al., 2023) (CL-13b), to provide a more comprehensive analysis across both open- and closed-source LLMs. Other variants of LLMut follow the same notation, e.g., LLMut (GPT-3.5) and LLMut (CL-13b). We compare these LLM-based mutant generation approaches with four widely-studied traditional approaches (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)), with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our primary experiments required the generation of more than 701,400 mutants.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutants generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutant rate, equivalent mutant rate, and efficiency). Additionally, we study how different prompt engineering strategies affect the effectiveness of the task. We also explore the potential of using LLMs for generating high-order mutants. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
@@
Our results reveal that, compared to existing approaches, LLMs generate more diverse mutants that are behaviorally closer to real bugs. This results in a higher fault detection rate compared to traditional approaches. For instance, LLM-based approaches achieve 1.75 improvement in real bug detection rate, with an increase of 32.99 percentage points. Specifically, the mutants generated by LLMut (DS-671b) achieve fault detection rates of 91.1%, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively. In terms of coupling rate, LLMorpheus (DS-671b) performs best, with 52% of its mutants coupled with real bugs. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutant rate, and equivalent mutant rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutants, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutant rate of 7.8%, and an equivalent mutant rate of 3.1%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that LLMut (DS-236b) and LLMut (DS-671b) exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutants. Moreover, the distributions of AST edit distance of mutants compared to the original code also show that LLMs generate more balanced and expressive mutants. For instance, almost 70% of the mutants of Major have an AST edit distance of one, whereas for the LLM-based approach, BugFarm (DS-671b), only 38% of its mutants have the same distance, indicating that LLM-based approaches tend to conduct more complex code transformations.
@@
We also categorise the mutants generated by GPT models that fail to compile, the most common issue with learning-based approaches generated mutants, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutants generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutants toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
@@
We analyze the surviving mutants of each approach, which are important to strengthen the test case. For one surviving non-equivalent mutant, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection. We analyze the distribution of surviving mutants across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutants, which consequently remain surviving, providing more valuable guidance for enhancing unit tests.
@@
We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our study not only covers other existing LLM-based mutant generation approaches, but also introduces a novel, effective prompt template. Our findings indicate that LLMs excel in generating diverse mutants that closely mimic real bugs.
We analyze the error types of non-compilable mutants and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutants.
We construct a high-quality dataset of Java mutants, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
@@ 2. Study Design
In this section, we introduce the design and conduct of our study.
@@ 2.1. Overview and Research Questions
This paper aims to evaluate existing mutant generation approaches by answering the following research questions.
@@
RQ1 (Effectiveness) How well do mutants generated by LLMs mimic real bugs compared to existing methods? This research question investigates the effectiveness of LLM-generated mutants in representing real bugs. Specifically, it evaluates whether these mutants exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity. This RQ is divided into the following sub-RQs:
@@
RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutants?
@@
RQ1.2 (Coupling) To what extent do LLM-generated mutants couple with real bugs?
RQ1.3 (Diversity) How diverse are LLM-generated mutants compared to traditional approaches?
@@
RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity? This research question explores the validity of LLM-generated mutants compared to traditional approaches. Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutants. By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutants that contribute to practical mutation testing. This RQ is divided into the following sub-RQs:
@@
RQ2.1 (Compilability) What percentage of generated mutants are compilable across different approaches?
@@
RQ2.2 (Duplication) What percentage of the generated mutants are duplicates of the original code across different approaches?
@@
RQ2.3 (Equivalence) What percentage of generated mutants are equivalent across different approaches?
@@
RQ3 (Efficiency) What are the general costs of different mutant generation approaches? This research question investigates the general costs of different mutant generation approaches. We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutant generation approaches.
@@
RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutants of different approaches? This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutants generated by different approaches. Understanding these errors can guide mutant generation approaches to enhance their practical usability.
@@
RQ5 (Surviving Mutants) What are the distributions of surviving mutants across different approaches? This RQ aims to investigate the underlying reasons why mutants remain surviving, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests. By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutants and their potential value for guiding test suite improvement.
2.2. Studied Mutant Generation Approaches
@@
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutant generation approaches are for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. The details of the settings for each generation approach are illustrated in Section 2.5. Table 1 summarizes the mutant generation approaches covered by this study.
@@
| BERT | Mutant generation with the pre-trained model CodeBERT. | |
@@
| BF (DS-671b) | The existing LLM-based approach BugFarm with DeepSeek-V3-671b. | |
| LLMph (GPT-4o) | The existing LLM-based approach LLMorpheus with GPT-4o. | |
| LLMph (DS-671b) | The existing LLM-based approach LLMorpheus with DeepSeek-V3-671b. | |
@@
| LLMut (DS-671b) | Our approach with DeepSeek-V3-671b, whose base model is DeepSeek, released in 2024/12. |